OpenStudy (anonymous):

Problem for Mathletes: Consider the positive integers that such that its digits are either \(1, 2,\) or \( 3\). Define a function \( f \) such that \( f(N) = \) the number of positive integers consisting of only \(1, 2 ,\) and \(3\) whose digits sum to \(N\). find \( f(10) \)

OpenStudy (anonymous):

cool. trying to understand what the problem wants

OpenStudy (mr.math):

NO HINTS PLEASE!

OpenStudy (zarkon):

274

OpenStudy (mr.math):

Not cool, zarkon! Don't explain your answer yet!

OpenStudy (mr.math):

This problem is giving me a headache, I'm still missing some cases.

OpenStudy (mr.math):

Especially the 7 digits case!

OpenStudy (mr.math):

I was able to get 194 only :(

OpenStudy (mr.math):

\[1+9+42+77+30+31+4\] @Zarkon, if you could tell us how you worked it out, that would be great. Thank you!

OpenStudy (mr.math):

Lets discuss it together. The maximum number of possible digits is 10, obviously: # of digits # of combinations 10: 1,1,1,1,1,1,1,1,1,1 1 9: 1,1,1,1,1,1,1,1,2 9 5: 3,3,2,1,1 OR 2,2,2,2,2 30+1 4: 3,3,3,1 4 3: - - These are probably the easy cases. I need to count for the 8, 7 and 6 digits.

OpenStudy (zarkon):

use a generating function

OpenStudy (anonymous):

Or you can use recurrence relation.

OpenStudy (asnaseer):

I used this method: 10x1 = 10 8x1 + 1x2 = C(9,1) = 9 7x1 + 1x3 = C(8,1) = 8 6x1 + 2x2 = C(8,2) = 28 5x1 + 1x3 + 1x2 = C(7,2) * C(2,1) = 21 * 2 = 42 4x1 + 2x3 = C(6,2) = 15 4x1 + 3x2 = C(7,3) = 35 3x1 + 1x3 + 2x2 = C(6,3) * C(3,1) = 20 * 3 = 60 2x1 + 2x3 + 1x2 = C(5,3) * C(3,1) = 10 * 3 = 30 2x1 + 4x2 = C(6,4) = 15 1x1 + 3x3 = C(4,3) = 4 1x1 + 1x3 + 3x2 = C(5,4) * C(4,1) = 5 * 4 = 20 2x3 + 2x2 = C(4,2) = 6 5x2 = 1 Total = 274